A car is at rest at the top of a driveway that has a slope of 14 degrees. If the brake of the car is released with the car in neutral gear find, a) the acceleration of the car down the drive and b) the time it takes for the car to reach the street, 8.0m away
Answer #1

Find the car's acceleration down the slope from

F = net force = weight component down slope

F=mg sin (theta)

a = F/m

a=mg sin (theta)/m

a=9.8(sin(14)=2.37m/s2

and s=(1/2)at^2

The displacement s=8.0m
8=(0.5)(2.37)t^2

t=2.59sec